OpenStudy (anonymous):

**WILL MEDAL** Which equation shows the substitution method being used to solve the system of linear equations? x + y = 6 x = y + 2 A. x + (y + 2) = 6 B. (y + 2) + y = 6 C. x + y = y + 2 D. x = (x-6) + 2

OpenStudy (anonymous):

@KamiBug I think its A

OpenStudy (kamibug):

Take the value of x, (y +2), and plug that in for x in the first equation. :) (y + 2) + y = 6 A is incorrect b/c it's substituting the value of x in for y instead of x.

OpenStudy (sachin00007):

the answer is B

OpenStudy (anonymous):

Thank u @KamiBug

OpenStudy (sachin00007):

it is because if you substitute x=y+2 into equation 1 you will get (y+2)+y=6

OpenStudy (anonymous):

Ok thx